6 years ago, Nilda was 5 times as old as Riza. in 5 years, nilda will be 3 times as old as riza. what is the present age of riza?
Answer by stanbon(75887)

6 years ago, Nilda was 5 times as old as Riza. in 5 years, Nilda will be 3 times as old as Riza. what is the present age of Riza?
-----------
Equations:
(n-6) = 5(r-6)
(n+5) = 3(r+5)
---
Rearrange:
n = 5r-30+6
n = 3r+15-5
--------------------
Substitute for n and solve for "r":
5r-24 = 3r+10
2r = 34
r = 17 (Riza's age now)
---
solve for "n":
n = 5r-24
n = 5*17-24
n = 85-24
n = 61 (Nilda's age now)
===========================
Cheers,
Stan H.
=============